OpenStudy (anonymous):

solve using distributive property: 7-(4h+10)=17

OpenStudy (anonymous):

To distribute multiply with -1..

OpenStudy (anonymous):

So you would get..\[7-4h-10=17\]

OpenStudy (anonymous):

do we lose the ()? or would it be 7 (-4h-10) = 17? trying to help my homeschooled daughter, but it has been awhile since i did this!

OpenStudy (anonymous):

Yeah. When multiplied the brackets are removed..

OpenStudy (anonymous):

So -5 is the answer?

OpenStudy (anonymous):

Yeah.. Well Done. ;)

OpenStudy (anonymous):

You could check..\[7-(4h+10)=17\]\[7-[4(-5)+10]=17\]\[7-(-20+10)=17\]\[7-(-10)=17\]\[7+10=17\]\[17=17\]

OpenStudy (anonymous):

Thanks so much, sure we will back!